OpenStudy (anonymous):

find the slope of a line -4y - 5x - 2 = 0

OpenStudy (unklerhaukus):

-4y-5x-2=0 -4y=5x+2 y=-5/4x -1/2 the equation is now of the form y=mx+b where m is the gradient and b the y-intercept

OpenStudy (anonymous):

- 5/4 ?

OpenStudy (unklerhaukus):

that is the slope

OpenStudy (chriss):

-4y-5x=2 -4y=5x+2 y=-5/4x-1/2 slope = -5/4

OpenStudy (anonymous):

thanks :)